Three of the primary components of air are carbon dioxide, nitrogen, and oxygen. In a sample containin a mixture of only these gases at exactly 1 atm, the partial pressures of carbon dioxide and nitrogen are given as


P

C

O

2

=

0.285

P

CO

2


=0.285


torr and


P

N

2

=

593.525

P

N

2


=593.525


torr. What is the partial pressure of oxygen?


1
Expert's answer
2022-01-30T07:58:17-0500

Partial pressure of CO2 = p = 0.285 torr = 0.000375 atm

Partial pressure of N2

= p'

593.525 torr = 593.525 / 760 atm

= 0.781 atm .


Partial pressure of O2 = p" .


Sum of all partial pressure = 1 atm .


P + p' + p" = 1 atm


P" = 1- p - p' = 1 - 0.000375 - 0.781 =

0.2186 atm